Paul got an 8% bump up in sales this week.Last week he made $10,000 in sales.How much extra money did he make this week
Answer by Theo(13342)

last week he made 10,000 in sales.

this week he made 8% more.

8% of 10,000 is equal to .08 * 10,000 which is equal to 8 * 100 = 800 dollars.

the bump up is an additional 800 dollars in sales.

800 / 10,000 = 8/100 = .08 = 8%.

alternatively you could solve as follows:

10,000 + .08 * 10,000 = x
x represents the new amount of sales.
factor out the 10,000 and you get 10,000 * (1 + .08) = x
simplify to get 1.08 * 10,000 = x
solve for x to get x = 10,800
10,800 minus 10,000 equals an additional 800 dollar in sales.